OpenStudy (anonymous):

What is the solution to the linear equation? 2/5 + p = 4/5 + 3/5p a. p = 1 b. p = 2 c. p = 8 d. p = 10

OpenStudy (igreen):

\(\dfrac{2}{5} + p = \dfrac{4}{5} + \dfrac{3}{5}p\) Subtract \(\dfrac{2}{5}\) to both sides..what's \(\dfrac{4}{5} - \dfrac{2}{5}\)? @justsmile531

OpenStudy (anonymous):

P=2

OpenStudy (anonymous):

4-2 is 2

OpenStudy (anonymous):

Look your ending equation is 2/5p divided by 4/5 and then you divide that getting p=2

OpenStudy (anonymous):

I was wrong because I forgot to simplify it sorry

OpenStudy (anonymous):

So what is the answer? 2 ?

OpenStudy (igreen):

No, it's 1.

OpenStudy (igreen):

@justsmile531

OpenStudy (anonymous):

Ok , thank you